What is the function y-3=2/3 (x-4) in the equivalent form most appropriate for identifying the slope and y-intercept of the function?
Answer by lwsshak3(11628)

What is the function y-3=2/3 (x-4) in the equivalent form most appropriate for identifying the slope and y-intercept of the function?
***
rewrite equation in standard form for a line: y=mx-b, m=slope, b=y-intercept
y-3=2/3(x-4)
3y-9=2x-8
3y=2x+1
y=2x/3+1/3
slope=2/3
y-intercept=1/3